# Efficient Long-duration Talking Video Synthesis with Linear Diffusion Transformer under Multimodal Guidance

### v2

@@ 3.1. Efficient Spatial-Temporal Transformer with Compression Autoencoder
Figure 3. Multimodal fusion schemes: (a) Direct Fusion injects conditions via cross-attention modules; (b) Siamese Fusion uses parallel transformer for feature guidance; (c) Symbiotic Fusion achieves fusion through input concatenation and self-attention. The backbone architecture (left-side blocks) remains consistent across all approaches.
@@
Table 1. Comparative analysis of three fusion schemes across three key dimensions: integration methods, condition states, and parameters sharing implementation
3.2. Analysis of Multimodal Fusion Schemes
@@
Integration methods: The choice between self-attention and cross-attention involves trade-offs in computational efficiency and optimization. Cross-attention, with smaller interaction matrices, is computationally lighter and easier to optimize, especially in limited data scenarios, as demonstrated in early text-to-image models (Rombach et al., 2022; Nichol et al., 2021). In contrast, self-attention, which requires larger matrices for multimodal token alignment, is more challenging to optimize but achieves higher embedding quality with sufficient data (Wang et al., 2024c). Alternatively, addition, which sums hidden states, provides a simpler and more computationally efficient approach, particularly useful when conditional inputs aim to highlight specific local features in the latent space.
@@
Condition States: The static versus dynamic nature of conditions reflects their constraint strength. Dynamic states, as in Siamese and Symbiotic Fusion, enable hierarchical refinement of conditional features during training, aligning with findings from ReferenceNet (Hu, 2024) and ControlNet (Zhang et al., 2023b) that multi-level conditioning enhances generation controllability. This flexibility improves feature alignment and model performance.
@@
Parameters Sharing: Siamese Fusion uses separate network parameters for conditions and latent noise, while Symbiotic Fusion shares the same parameters. Direct Attention, relying solely on cross-attention, also does not share parameters.
@@
Modality Adaptation: The adaptability of fusion schemes to different modalities is key to their success in multimodal tasks. Direct Fusion, though offering weak alignment, is highly cost-effective for modalities with significant differences, such as text and image, making it ideal for resource-constrained scenarios. Siamese Fusion offers enhanced design flexibility by allowing adjustable modality alignment. This is exemplified in models such as ControlNet (Zhang et al., 2023b), which employs visual cue-level alignment, and AnimateAnyone (Hu, 2024), which utilizes pixel-level alignment. The strength of alignment can thus be precisely tailored to meet the specific requirements of the task at hand. Symbiotic Fusion enhances parameter efficiency compared to Siamese Fusion, achieving strong alignment with fewer parameters. This makes it particularly effective for large-scale applications, such as AnimateAnyone (Hu, 2024), where shared parameters ensure robust and efficient multimodal integration. The comparative analysis of the three fusion approaches are summarized in Tab. 1
3.3. Guidance on Reference Portrait and Audio
Given the reference image, we require that the character of the generated video is consistent with the reference. Since the image modality is close to the video, Symbiotic Fusion proves appropriate. Specifically, we first encode the reference image and obtain the portrait feature $\bm{P}\in\mathbb{R}^{P\times d}$ in the latent space. Then we temporally repeat the portrait feature $\bm{P}$ with $F$ frames and concatenate it with the latent feature $\bm{Z_{s}}\in\mathbb{R}^{F\times P\times d}$ along the spatial dimension, which can be formulated as: $\bm{Z}=\text{concat}\left(\,\bm{Z_{s}},\,\tilde{\bm{P}}\,\right)$ denotes the temporally extended portrait feature. This operation yields a concatenated latent feature $\bm{Z}\in\mathbb{R}^{F\times 2P\times d}$. By leveraging shared self-attention in the fused feature space, this method maintains consistent character representation across frames. In contrast, Direct Fusion uses shallow portrait features, leading to misalignment in deeper layers. While Siamese Fusion mitigates this by maintaining parallel transformer branches, it increases computational complexity without significant advantages over Symbiotic Fusion.
@@ 3.3. Guidance on Reference Portrait and Audio
As for the audio, we require it to reasonably drive the animation, especially the lip motion (Zhang et al., 2023a). However, certain unconscious actions (e.g. blinking and head movements) are almost independent of the audio yet serve to diversity. Further considering that the modality of audio is quite different from the video, using Siamese Fusion and Symbiotic Fusion often causes ambiguity in the backbone learning, damaging the diversity. In a nutshell, Direct Fusion is concise yet effective for fusing the input audio. In addition, for the audio representation, to accommodate the tokenization mechanism of the transformer, we follow (Xu et al., 2024b) and extract richer semantic information from different layers of the wav2vec encoder (Schneider et al., 2019). Through nonlinear mapping, we obtain audio tokens $\bm{C}\in\mathbb{R}^{F\times T\times d^{\prime}}$ enriched with semantic content, which are aligned with latent features in the same feature space. Temporal-level audio representations are then synchronized with video sequences, resulting in the semantic audio representation. The experiments in Sec. 4.4 verify the efficacy of our mechanism in feature fusion.
Figure 4. The illustration of the long-duration generation.
3.4. Long-Duration Generation
Unlike the previous frame-by-frame generation approach (Shen et al., 2023; Stypułkowski et al., 2024), our end-to-end diffusion transformer generates a video clip in a single sampling process. In diffusion-based video generation tasks, ensuring proper temporal alignment is critical for maintaining coherence and consistency across generated video sequences. Additionally, due to the enhanced realism and dynamism of the talking videos we generate, direct transitions between consecutive clips are often not feasible. To address this, we introduce a memory bank that stores information from previously generated clips. This data undergoes patchification and temporal transpose operations, transforming it into $\bm{M}\in\mathbb{R}^{P\times f\times d}$, where $f$ represents the number of frames selected from the prior clips. Leveraging the scalability of the Transformer architecture, the memory bank directly provides prior information through Symbiotic Fusion with $\bm{Z_{t}}$. However, the inherent distribution discrepancy between training (real data) and inference (synthetic data) poses a fundamental challenge for diffusion models, with error accumulation becoming increasingly severe as video length grows.
To mitigate this latent risk, we develop a noise-regularized training strategy. Specifically, we first establish that diffusion models inherently retain residual noise even after finite-step denoising due to the discrete approximation of continuous noise schedules, which progressively accumulates as the video sequence extends. Building on the insight, we introduce a bounded stochastic noiser that perturbs the memory bank contents with controlled Gaussian noise $\epsilon\sim\mathcal{N}(0,{\sigma_{mini}}^{2}\mathbf{I})$ (see Fig. 4) into the memory bank during training. Inspired by Hallo2 (Cui et al., 2024), we further apply a patch-drop data augmentation strategy to the memory bank. Unlike Hallo2 which protects face patches from dropping, we employ completely random dropping to enhance the model’s robustness to the memory bank. This deliberate corruption serves two purposes: (1) it faithfully simulates inference-time distribution shifts, and (2) it preserves crucial temporal dependencies. The resulting framework demonstrates significantly improved robustness against error propagation in long-duration generation tasks.
@@ 3.4. Long-Duration Generation
Figure 5. The qualitative comparisons with other cutting-edge methods on the HDTF dataset. Our method achieves better audio-animation alignment (e.g. lip motions) and produces expressive results.
4. Experiment
4.1. Experimental Setups
Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF (Zhang et al., 2021) and CelebV-HQ (Zhu et al., 2022). These datasets include videos that consist of individuals of various ages, ethnicities, and genders, captured in half-body or close-up shots against varied indoor and outdoor backgrounds. To ensure high-quality training data, we adopt the preprocess scheme used in (Xu et al., 2024b). We first exclude videos with scene changes, significant camera movements, excessive facial motion, and side-facing shots. Then we refine the remaining videos and retain only single-person speaking videos with strong audio-lip correspondence. After refinement, we obtain 356 clips with 13.95 hours from the HDTF, and 24438 clips with 44.91 hours from the CelebV-HQ dataset. During training, we extract 16-frame video clips from these datasets using a specific sampling interval, each frame is resized to $256\times 256$ resolution.
Evaluation Metrics The evaluation metrics utilized in the portrait image animation approach include Fréchet Inception Distance (FID), Fréchet Video Distance (FVD), Structural Similarity Index Measure (SSIM) and Expression-FID (E-FID). Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs (Karras et al., 2019). SSIM indices measure the structural similarity between ground truth videos and generated videos. E-FID leverages Inception features to assess image authenticity, providing a more precise fidelity measure. It extracts facial expression parameters via a reconstruction method (Deng et al., 2019) and computes FID to quantify the gap between generated and ground truth expressions.
Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) (Ma et al., 2023), SadTalker (138M) (Zhang et al., 2023a), AniPortrait (2.8B) (Wei et al., 2024), Hallo (2.4B) (Xu et al., 2024b), Hallo (2.4B) (Cui et al., 2024) and EchoMimic (2.6B) (Chen et al., 2024). The evaluation is performed on the HDTF and CelebV-HQ Dataset, utilizing a training and testing split where $90\%$ of the identity data was allocated for training purposes.
@@ 4.1. Experimental Setups
Figure 6. The qualitative comparisons with the existing portrait image animation approaches on the CelebV-HQ dataset. Our method achieves better portrait consistency.
@@
DreamTalk 105.434 529.510 0.406 3.466
SadTalker 102.371 362.634 0.414 2.168
AniPortrait 78.284 331.117 0.623 2.903
EchoMimic 55.791 299.343 0.639 2.493
Hallo 45.754 285.988 0.698 2.235
Hallo2 44.343 241.628 0.657 2.017
@@
Table 2. The quantitative comparisons with the existed portrait image animation approaches on the HDTF dataset.
Implementation details Our approach, in contrast to previous methods (Tian et al., 2024; Xu et al., 2024b; Wei et al., 2024) that rely on a two-stage training process, adopts an end-to-end training strategy. We set up models with two sizes: the Base model, which is suitable for simple fine-tuning, and the Large model, which demonstrates better generalization ability.
4.2. Quantitative Results
The comprehensive quantitative comparisons of HDTF and CelebV datasets are listed in Tab. 2 and Tab. 3, respectively. The results show that our method significantly outperforms other state-of-the-art methods in almost all metrics, achieving the best portrait consistency and audio-animation alignment. Notably, our approach achieves the lowest FID (31.761) and FVD (219.119) on the HDTF dataset, as well as the lowest FID (7.610) and FVD (33.175) on the CelebV-HQ dataset. These results show that our method generates both realistic and temporally consistent talking-head videos. For a fair comparison, each method generates the first video clip for every validation video, which is then resized to the same resolution for evaluation.
4.3. Qualitative Results
@@
DreamTalk 64.297 293.905 0.448 5.915
SadTalker 42.017 77.802 0.469 3.372
AniPortrait 33.408 107.236 0.578 3.722
EchoMimic 12.932 59.944 0.567 3.671
Hallo 8.504 46.431 0.580 3.280
Hallo2 8.734 41.213 0.543 3.128
@@
Table 3. The quantitative comparisons with the existed portrait image animation approaches on the CelebV-HQ dataset.
@@
Table 4. Experimental results of three multimodal fusion schemes, guided by reference portraits, on the HDTF validation dataset, employing the Letstalk-B model.
@@
Table 5. Experimental results of three multimodal fusion schemes, driven by audio inputs, on the HDTF validation dataset, employing the Letstalk-B model.
4.4. Ablation Study
@@
Analysis on the Reference Portrait Fusion Portrait consistency is crucial to the portrait image animation task and highly relevant to the portrait fusion scheme used. We evaluate the synthesis quality of portrait fusion using three schemes, and the quantitative results are shown in Tab. 4. Symbiotic Fusion almost surpasses the other two schemes, realizing better portrait consistency and video quality. Besides, while Siamese Fusion produces comparable results, it introduces additional parameters (i.e. siamese transformer) and complicates the learning task.
@@
Analysis on the Driven Audio Fusion The way we fuse audio affects both audio-animation alignment and diversity. Audio-animation alignment refers to specific visual cues (Zhang et al., 2023a) (e.g. facial muscle and lip motions), whereas the diversity is related to the speech-agnostic motions (e.g. blinking and neck moving) that often appear as unconscious actions. We observe that diversity mainly comes from the transformer backbone during denoising. Thus, we aim to fuse the audio to achieve alignment while preserving the backbone’s inherent diversity. Quantitative results in Tab. 5 indicate that Direct Fusion outperforms the other schemes in establishing the correspondence between audio and animation.
Method FID $\downarrow$ FVD $\downarrow$
Baseline model with Memory Bank 52.243 276.16
+ Regularized Training (w/o patch drop) 32.746 264.709
+ Regularized Training (w patch drop) 34.821 251.684
Table 6. Experimental results of long video generation. Experimental results comparing three approaches for 64-second video generation (100 clips): (1) memory bank baseline, (2) regularized training, and (3) patch drop augmentation.
4.5. Long-Duration Generation
To validate the effectiveness of our proposed approach for long video generation, we conducted experiments to analyze the quality of 100 consecutive video clips (about 64 second) generated under three settings: (1) the baseline model (LetsTalk-L) with with the memory bank, (2) the method augmented with regularized training strategy, and (3) the method further enhanced with the patch drop augmentation. Our results in Tab. 6 demonstrate that the memory bank significantly improves temporal consistency, while the regularized training strategy further mitigates degradation in later clips, leading to higher overall video quality. The results demonstrate that regularized training strategy primarily enhances single-frame quality, significantly reducing FID from 52.243 to 32.746, while maintaining temporal consistency (FVD 264.709). On the other hand, patch drop augmentation further improves long-term motion coherence, achieving the best FVD (251.684) with only a slight FID increase (34.821), confirming its complementary role in stabilizing long-duration generation.
5. Conclusion
We introduced LetsTalk, an efficient linear diffusion transformer for high-fidelity talking head synthesis, tackling core challenges in multimodal fusion and long-term temporal coherence. Our approach integrates a deep compression autoencoder for compact latent representation, alongside a systematic exploration of fusion strategies—yielding an optimal framework with Symbiotic Fusion for portrait consistency and Direct Fusion for audio-video synchronization. The memory bank mechanism and noise-regularized training further mitigate error accumulation in extended sequences. Extensive experiments demonstrate LetsTalk’s state-of-the-art performance, generating lifelike, temporally coherent animations with enhanced diversity, all while maintaining superior efficiency through its optimized architecture.
Appendix A Preliminaries
(2) $\begin{split}\mathcal{L}(\theta)=&\sum_{t}\mathcal{D}_{\text{KL}}(q^{*}(x_{t-1}|x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}|x_{t}))\\ &-p(x_{0}|x_{1}).\end{split}$
(3) $\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x)-\epsilon(x)\|_{2}^{2}.$
Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework. DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches. This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently. Different from ViT, DiT incorporates a variational autoencoder (VAE) (Kingma, 2013) in the framework, thus runs in the latent space, rather than pixel space. Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.
Model Patch size Layer num N Hidden size D Paramaters
LetsTalk-B 1 12 768  [^2] M
LetsTalk-L 1 24 1024 773
Table 7. Details of LetsTalk models. We follow ViT and DiT model configurations for different model sizes.
Appendix B Network Architecture Details
B.1. Backbone
In LetsTalk, we borrow the Variational AutoEncoder (VAE) (Kingma, 2013) from Stable Diffusion (Rombach et al., 2022). Given the reference portrait image and noisy video, we first encode them using the encoder in VAE. Then we patchify and flatten them along the spatial dimension. Finally, we use an MLP to project them into the latent space. The embeddings in the latent space will be fed into our transformer backbone. Specifically, our LetsTalk model is constructed using a series of N transformer blocks, each with a hidden dimension of D and H multi-head attention heads. Inspired by ViT and DiT, we explore two configurations of LetsTalk with varying numbers of parameters, while maintaining a fixed patch size of 2, as detailed in  Tab. 7. Via the encoder in VAE, the downsampling factor is $8$, the input reference portrait image $\bm{I}_{\text{p}}\in\mathbb{R}^{256\times 256\times 3}$ and the noisy video $\bm{V}\in\mathbb{R}^{16\times 256\times 256\times 3}$ are encoded as a portrait feature $\bm{P}\in\mathbb{R}^{32\times 32\times 4}$ and a video feature $\bm{F}\in\mathbb{R}^{16\times 32\times 32\times 4}$. Next, the two features are partitioned according to $2\times 2$ patch size and flattened as $\bm{P}\in\mathbb{R}^{256\times 16}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 16}$. Take our base model (i.e. LetsTalk-Base) as an example, we use an MLP to project both features into the latent space as $\bm{P}\in\mathbb{R}^{256\times 768}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 768}$.
B.2. Adaptive Integration
Following VDT(Lu et al., 2023), we insert an adaptive full-connected layer between attention layers:
(4) $\bm{Z}_{s}=\bm{Z_{s}}+\alpha\cdot\text{FC}(\text{Temp-Attn}(\bm{Z_{t}}))$
B.3. Audio Encoder
We resample the audio signal to a 16 KHz sample rate. For audio feature extraction, we use wav2vec (Schneider et al., 2019), employing embeddings from the last 12 layers to capture a richer range of semantic information across different levels of the audio representation. To ensure frame-by-frame alignment between audio and animation, we match the audio features with corresponding frames of the video clips. Additionally, we use a sliding window (multiple audio frames) for each video frame to maintain audio continuity and preserve more accurate audio information. To extract audio representations suitable for transformer tokenization, we designed an audio projection module. This module first flattens the extracted audio features and then selects a fixed number of context tokens (32 in our experiments). Simultaneously, the audio representations are mapped to the same dimensional feature space as the latent noise.
B.4. Siamese Transformer
The Siamese Fusion scheme has been adopted in AnimateAnyone (Hu, 2024), EMO (Tian et al., 2024), Hallo (Xu et al., 2024b), and so on. The Siamese Fusion scheme aims to update a backbone-like network simultaneously and extract the features to fuse it into the backbone block-wise. Since these methods follow Stable Diffusion and use a U-Net to predict the noise for denoising, they adopt a similar U-Net as the siamese network. Considering LetsTalk as a Diffusion Transformer, we introduce a ViT (Dosovitskiy, 2020) pre-trained on ImageNet to achieve the Siamese Fusion scheme for the reference portrait image. For the input audio, we propose a transformer whose blocks consist of a two-layer MLP to implement the Siamese Fusion scheme.
@@ B.4. Siamese Transformer
Figure 7. Portrait animation results without audio guidance for three fusion schemes. Siamese and Symbiotic fusion schemes with high fusion compactness limit the diversity of video generation, while Direct fusion preserves this diversity.
B.5. Impact of Audio-guidance on Diversity

### v3

@@ III-C Memory Bank and Noise-Regularized Training for Long-Term Consistency
Figure 4: Multimodal fusion schemes: (a) Direct Fusion injects conditions via cross-attention modules; (b) Siamese Fusion uses parallel transformer for feature guidance; (c) Symbiotic Fusion achieves fusion through input concatenation and self-attention. The backbone architecture (left-side blocks) remains consistent across all approaches.
@@
TABLE I: Comparative analysis of three fusion schemes across three key dimensions: integration methods, condition states, and parameters sharing implementation
III-D Analysis of Multimodal Fusion Schemes
@@
Integration methods: The choice between self-attention and cross-attention involves trade-offs in computational efficiency and optimization. Cross-attention, with smaller interaction matrices, is computationally lighter and easier to optimize, especially in limited data scenarios, as demonstrated in early text-to-image models [13, 52]. In contrast, self-attention, which requires larger matrices for multimodal token alignment, is more challenging to optimize but achieves higher embedding quality with sufficient data [53]. Alternatively, addition, which sums hidden states, provides a simpler and more computationally efficient approach, particularly useful when conditional inputs aim to highlight specific local features in the latent space.
@@
Condition States: The static versus dynamic nature of conditions reflects their constraint strength. Dynamic states, as in Siamese and Symbiotic Fusion, enable hierarchical refinement of conditional features during training, aligning with findings from ReferenceNet [17] and ControlNet [14] that multi-level conditioning enhances generation controllability. This flexibility improves feature alignment and model performance.
@@
Parameter Sharing: Siamese Fusion uses separate network parameters for conditions and latent noise, while Symbiotic Fusion shares the same parameters. Direct Fusion, relying solely on cross-attention, also does not share parameters. Parameter sharing in Symbiotic Fusion is particularly advantageous for large-scale training, improving efficiency and scalability.
@@
Modality Adaptation. The adaptability of fusion schemes to different modalities is key to their success in multimodal tasks. Direct Fusion, though offering weak alignment, is highly cost-effective for modalities with significant differences, such as text and image, making it ideal for resource-constrained scenarios. Siamese Fusion offers enhanced design flexibility by allowing adjustable modality alignment. This is exemplified in models such as ControlNet [14], which employs visual cue-level alignment, and AnimateAnyone [17], which utilizes pixel-level alignment. The strength of alignment can thus be precisely tailored to meet the specific requirements of the task at hand. Symbiotic Fusion enhances parameter efficiency compared to Siamese Fusion, achieving strong alignment with fewer parameters. This makes it particularly effective for large-scale applications, such as AnimateAnyone [17], where shared parameters ensure robust and efficient multimodal integration. The comparative analysis of the three fusion approaches is summarized in Table I.
In summary, we adopt Symbiotic Fusion for portrait features to ensure strong visual consistency, and Direct Fusion for audio features to maintain motion diversity and effective audio-visual alignment.
III-E Guidance on Reference Portrait and Audio
[…]
@@ III-E Guidance on Reference Portrait and Audio
As for the audio, we require it to reasonably drive the animation, especially the lip motion [1]. However, certain unconscious actions (e.g. blinking and head movements) are almost independent of the audio yet serve to increase diversity. Considering that the modality of audio is quite different from the video, using Siamese Fusion and Symbiotic Fusion often causes ambiguity in the backbone learning, damaging the diversity. In a nutshell, Direct Fusion is concise yet effective for fusing the input audio.
For the audio representation, to accommodate the tokenization mechanism of the transformer, we follow [18] and extract richer semantic information from different layers of the wav2vec encoder [54]. Through nonlinear mapping, we obtain audio tokens $\bm{C}\in\mathbb{R}^{F\times T\times d^{\prime}}$ enriched with semantic content, which are aligned with latent features in the same feature space. Temporal-level audio representations are then synchronized with video sequences, resulting in the semantic audio representation. The experiments in Section IV-F verify the efficacy of our mechanism in feature fusion.
IV Experiment
@@
Figure 5: The qualitative comparisons with other cutting-edge methods on the HDTF dataset. Our method achieves better audio-animation alignment (e.g. lip motions) and produces expressive results.
@@
Figure 6: The qualitative comparisons with the existing portrait image animation approaches on the CelebV-HQ dataset. Our method achieves better portrait consistency.
IV-A Experimental Setups
Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF [55] and CelebV-HQ [56]. These datasets include videos that consist of individuals of various ages, ethnicities, and genders, captured in half-body or close-up shots against varied indoor and outdoor backgrounds. To ensure high-quality training data, we adopt the preprocess scheme used in [18]. We first exclude videos with scene changes, significant camera movements, excessive facial motion, and side-facing shots. Then we refine the remaining videos and retain only single-person speaking videos with strong audio-lip correspondence. After refinement, we obtain 356 clips with 13.95 hours from the HDTF, and 24438 clips with 44.91 hours from the CelebV-HQ dataset. During training, we extract 16-frame video clips from these datasets using a specific sampling interval, each frame is resized to $512\times 512$ resolution.
Evaluation Metrics The evaluation metrics utilized in the portrait image animation approach include Fréchet Inception Distance (FID), Fréchet Video Distance (FVD), Structural Similarity Index Measure (SSIM) and Expression-FID (E-FID). Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs [57]. SSIM indices measure the structural similarity between ground truth videos and generated videos. E-FID leverages Inception features to assess image authenticity, providing a more precise fidelity measure. It extracts facial expression parameters via a reconstruction method [58] and computes FID to quantify the gap between generated and ground truth expressions.
Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) [33], SadTalker (138M) [1], AniPortrait (2.8B) [2], Hallo (2.4B) [18], Hallo 2 (2.4B) [20] and EchoMimic (2.6B) [19]. The evaluation is performed on the HDTF and CelebV-HQ Dataset, utilizing a training and testing split where $90\%$ of the identity data was allocated for training purposes.
Figure 7: English audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
@@ IV-A Experimental Setups
DreamTalk [33] 105.434 529.510 0.406 3.466
SadTalker [1] 102.371 362.634 0.414 2.168
AniPortrait [2] 78.284 331.117 0.623 2.903
EchoMimic [19] 55.791 299.343 0.639 2.493
Hallo [18] 45.754 285.988 0.698 2.235
Hallo2 [20] 44.343 241.628 0.657 2.017
@@
TABLE II: The quantitative comparisons with the existed portrait image animation approaches on the HDTF dataset.
@@
DreamTalk [33] 64.297 293.905 0.448 5.915
SadTalker [1] 42.017 77.802 0.469 3.372
AniPortrait [2] 33.408 107.236 0.578 3.722
EchoMimic [19] 12.932 59.944 0.567 3.671
Hallo [18] 8.504 46.431 0.580 3.280
Hallo2 [20] 8.734 41.213 0.543 3.128
@@
TABLE III: The quantitative comparisons with the existed portrait image animation approaches on the CelebV-HQ dataset.
IV-B Additional Implementation Details
Unlike previous methods [36, 18, 2] that use a two-stage training process, our approach adopts an end-to-end training strategy. We provide two model sizes: a Base model for simple fine-tuning, and a Large model with better generalization.
Diffusion We follow the hyperparameters from ADM [12], using a linear variance schedule with $t_{\text{max}}=1000$, ranging from $1\times 10^{-4}$ to $2\times 10^{-2}$. We also use ADM’s parameterization of the covariance $\Sigma_{\theta}$ and their method for timestep embedding.
Training LetsTalk is initialized from a pre-trained Latte [15] model. We use gradient clipping for stability and the AdamW optimizer with a fixed learning rate of $1\times 10^{-4}$ and no weight decay, following Latte’s training setup. As in previous generative modeling work [29, 59], we maintain an exponential moving average (EMA) of model weights with a decay rate of $0.9999$, and report all results using the EMA model.
Inference At inference, the model takes a single reference portrait and driving audio as input, and generates a video sequence that animates the reference image according to the audio. The generated video has the same length and resolution as during training: $16$ frames at $512\times 512$ resolution.
IV-C Quantitative Results
We report comprehensive quantitative comparisons on the HDTF and CelebV-HQ datasets in Table II and Table III, respectively. As shown in the tables, our method achieves the best performance on most metrics. Specifically, on the HDTF dataset, LetsTalk-L achieves the lowest FID (31.761) and FVD (219.119), outperforming the previous best method Hallo2 by 28.3% and 9.3%, respectively. On CelebV-HQ, our method also achieves the lowest FID (7.610) and FVD (33.175). These results demonstrate that our approach generates more realistic and temporally consistent talking-head videos. All methods generate the first video clip per validation sample and are evaluated at the same resolution.
IV-D Qualitative Results
Figure 8: Chinese audio-driven results of LetsTalk. Note that each column corresponds to the same audio.
Method #Params Inference Time (ms/frame) FID $\downarrow$
SadTalker [1] 138M 23 93
EchoMimic [19] 2.2B 54 55
AniPortrait [2] 2.6B 74 82
Hallo [18] 2.4B 115 44
LetsTalk-B (Ours) 0.3B 27 46
LetsTalk-L (Ours) 0.7B 40 37
TABLE IV: Comparison of model size, inference time (ms per frame), and FID on the HDTF dataset.
IV-E Efficiency Comparison
To further demonstrate the practical value of our method, we compare model size, inference speed, and FID score with previous state-of-the-art methods. As shown in Table IV, LetsTalk-B achieves comparable or better visual quality with significantly fewer parameters and faster inference than recent diffusion-based approaches. The larger LetsTalk-L further improves FID while maintaining reasonable efficiency. These results highlight the strong trade-off between quality and efficiency achieved by LetsTalk, making it well-suited for real-world applications.
IV-F Ablation Study
@@
TABLE V: Experimental results of three multimodal fusion schemes, guided by reference portraits, on the HDTF validation dataset, employing the Letstalk-B model.
@@
TABLE VI: Experimental results of three multimodal fusion schemes, driven by audio inputs, on the HDTF validation dataset, employing the Letstalk-B model.
@@
Analysis on the Reference Portrait Fusion Portrait consistency is crucial to the portrait image animation task and highly relevant to the portrait fusion scheme used. We evaluate the synthesis quality of portrait fusion using three schemes, and the quantitative results are shown in Table V. Symbiotic Fusion almost surpasses the other two schemes, realizing better portrait consistency and video quality. Besides, while Siamese Fusion produces comparable results, it introduces additional parameters (i.e. siamese transformer) and complicates the learning task.
@@
Analysis on the Driven Audio Fusion The way we fuse audio affects both audio-animation alignment and diversity. Audio-animation alignment refers to specific visual cues [1] (e.g. facial muscle and lip motions), whereas the diversity is related to the speech-agnostic motions (e.g. blinking and neck moving) that often appear as unconscious actions. We observe that diversity mainly comes from the transformer backbone during denoising. Thus, we aim to fuse the audio to achieve alignment while preserving the backbone’s inherent diversity. Quantitative results in Table VI indicate that Direct Fusion outperforms the other schemes in establishing the correspondence between audio and animation.
@@
Figure 9: Portrait animation results without audio guidance for three fusion schemes. Siamese and Symbiotic fusion schemes with high fusion compactness limit the diversity of video generation, while Direct fusion preserves this diversity.